OpenStudy (anonymous):

You roll a pair of dice. What is the probability that both dice will show a multiple of 3? Answer choices: 2/3 4/9 1/3 1/9

OpenStudy (campbell_st):

well there are 36 possible outcomes when you roll 2 dice and multiples of 3 you can have 3, 3 or 3, 6 or 6, 3 or 6, 6 so you have 4/36 just simplify the fraction

OpenStudy (anonymous):

\[\left(\frac{2}{6}\right)^2=\frac{1}{9} \]